### v1

@@ 4.2.3 ISA-extended Sparse Kernel
We prototyped the functionality of the xDecimate instruction at the register-transfer level (RTL) inside the open-source RI5CY/CV32E40P processor  [^2] https://github.com/openhwgroup/cv32e40p used in Rossi et al. (2021). RI5CY implements the RV32IMC ISA with the addition of the Xpulpv2 extension, which includes hardware loops, load/store with post-increment, SIMD dot-product instructions, etc. We modified the processor adding an eXtension Functional Unit (XFU), whose microarchitecture is shown in Fig. 7, which encompasses three of the four stages of RI5CY’s pipeline: Instruction Decode (ID), Execute (EX), and Write-Back (WB). Differently from regular RISC-V loads, the xDecimate instruction employs an R-type encoding with two source registers rs1, rs2 and one destination register rd; all three registers are simultaneously read from the register file (RF, not shown in Fig. 7) in the ID stage, exploiting RI5CY’s 3-read port RF, which is also necessary for the Xpulpv2 extension. A lightweight decoder identifies which flavor of xDecimate (i.e., which sparsity format) is being used.
@@ 5 Results
5.1 Experimental Setup
[…]
5.2 Single Layers Benchmarking
5.3 End-to-end DNNs Benchmarking
@@ 6 Conclusions
We have introduced a set of efficient kernels N:M pruning on MCUs of the PULP family, and a lightweight extension of the XPulpV2 ISA, to further speed them up. On the platform described in Rossi et al. (2021), and targeting a ViT and a ResNet18, we have achieved end-to-end latency reductions of 1.81$\times$ and 3.21$\times$, with less than 0.5/1.5% accuracy drop on CIFAR10/CIFAR100 respectively.

### v2

@@ 4 Lightweight Kernels for N:M Sparse DNNs on MCUs
As outlined in Section 1, the novelty of our work lies in being the first, to the best of our knowledge, to propose a way to accelerate N:M sparsity on MCUs, either through SW or through an area-inexpensive HW extension, without relying on ad-hoc accelerators. The aim is to unlock a new set of Pareto-optimal trade-offs in the space of model accuracy vs. latency reduction vs. area cost, bridging the gap between the significant speedups offered by dedicated HW accelerators and classic MCUs without HW and SW extensions.
@@ 4.1.2 Software-only Sparse Kernel
Notably, although our implementation is partly target-specific, the same design choices performed for our kernels can be extended to support other MCUs, without relying on features like on-board DMA or the XPulpV2 ISA extension. For instance, ARM Cortex-M architectures leverage CMSIS-NN Lai et al. (2018) kernels and SIMD operations for dense convolutions, thus very similar kernel routines can be designed to add support for N:M sparsity.
@@ 4.2.3 ISA-extended Sparse Kernel
We prototyped the functionality of the xDecimate instruction at the register-transfer level (RTL) inside the open-source RI5CY/CV32E40P processor  [^2] https://github.com/openhwgroup/cv32e40p used in Rossi et al. (2021). The RTL is available open-source at https://github.com/eml-eda/cv32e40x_deci
mate
. RI5CY implements the RV32IMC ISA with the addition of the Xpulpv2 extension, which includes hardware loops, load/store with post-increment, SIMD dot-product instructions, etc. We selected a RISC-V processor for implementing our xDecimate instruction, leveraging the inherent extensibility of the ISA. However, in principle, our hardware extension could also be adapted to other architectures, such as ARM.
We modified the processor by adding an eXtension Functional Unit (XFU), whose microarchitecture is shown in Fig. 7, which encompasses three of the four stages of RI5CY’s pipeline: Instruction Decode (ID), Execute (EX), and Write-Back (WB). Differently from regular RISC-V loads, the xDecimate instruction employs an R-type encoding with two source registers rs1, rs2 and one destination register rd; all three registers are simultaneously read from the register file (RF, not shown in Fig. 7) in the ID stage, exploiting RI5CY’s 3-read port RF, which is also necessary for the Xpulpv2 extension. A lightweight decoder identifies which flavor of xDecimate (i.e., which sparsity format) is being used.
@@ 5 Results
5.1 Experimental Setup
[…]
After training, we quantize all networks to 8 bits using Brevitas Pappalardo (2023). All individual kernels and end-to-end ResNets have been deployed using the modified MATCH compiler. For ViTs, given the lack of support for attention layers in MATCH, we computed the latency layer-by-layer, using Deeploy Scherer et al. (2024), an alternative compiler for the same HW, for attention layers, and MATCH for sparse/dense feed-forward layers. The reported MACs/cycle values always refer to dense-equivalent operations.
5.2 Single Layers Benchmarking
5.3 End-to-end DNNs Benchmarking
While our benchmark DNNs are relatively small, compatibly with edge applications’ constraints, we underline that even greater speedups would be achieved on bigger models, given the combined effect of a more aggressive pruning applicable to them and the lower impact of overhead effects (e.g., external loop management) on bigger layers, in which most of the complexity resides in the internal convolutional kernel loops that we optimize.
5.4 SoTA Comparison
@@ 6 Conclusions
We have introduced a set of efficient kernels N:M pruning on MCUs of the PULP family, and a lightweight extension of the XPulpV2 ISA, to further speed them up. On the platform described in Rossi et al. (2021), and targeting a ViT and a ResNet18, we have achieved end-to-end latency reductions of 1.81$\times$ and 3.21$\times$, with less than 0.5/1.5% accuracy drop on CIFAR10/CIFAR100 respectively. Our future work will study the impact of variable sparsity patterns (e.g., per-layer or per-channel) on latency and accuracy, considering both the pure software kernels and the xDecimate-enhanced ones. Further, we will prototype our hardware extension on FPGA to enable an estimation of the energy savings achieved by our kernels, which can show further advantages in the reduced off-chip memory accesses.